OpenStudy (anonymous):

Use elimination to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) {12x+y=4-2x−2y=-6

OpenStudy (anonymous):

no, maybe \[12x+y=4\\ -2x-2y=-6\]

OpenStudy (anonymous):

how fast you want to do it?

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

we didn't do anything yet, i just wrote down what i thought the question was

OpenStudy (anonymous):

lol so you really must be confused !! lets go slow

OpenStudy (anonymous):

i wrote it wrong its 1/2x + y= 4 -2x-2y=-6

OpenStudy (anonymous):

oooh ok lets get rid of annoying fraction right away by multiplying everything in the first equation \[\frac{1}{2}x+y=4\] by \(4\)

OpenStudy (anonymous):

we are going to multiply each term by \(4\) and get \[2x+4y=16\] for the first one clear or no?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok now we line them up \[2x+4y=16\\ -2x-2y=-6\]

OpenStudy (anonymous):

add in a column, and a miracle will occur, since \(2x-2x=0\) you end up with \[2y=10\] when you add

OpenStudy (anonymous):

ok to that step or no?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then if \(2y=10\) is is pretty clear what \(y\) is yes?

OpenStudy (anonymous):

please say "yes" (i hope)

OpenStudy (anonymous):

5

OpenStudy (anonymous):

mm

OpenStudy (anonymous):

thanks